Top shellfish aquaculture product, produced in Newfound-land and Labrador, Nova Scotia, Prince Edward Island and Quebec. Fish filter the sea water between their shells and accumulate viral, bacterial, and parasitic pathogens in their tissues. Among the Fish collected from a certain Canadian coastline 80% are infected with a parasite. A biologist examines 120 randomly selected fish in this region. What is the probability that more than 86% of the fish in this sample are infected with the parasite?
Answer #1

Sampling distribution of p̂ is approximately normal if np >=10 and n (1-p) >= 10
n * p = 120 * 0.8 = 96
n * (1 - p ) = 120 * (1 - 0.8) = 24

Mean = \mu_{\hat{p}} = p = 0.8

Standard deviation = \sqrt{ p (1 - p) / n} = 0.036515

X ~ N ( µ = 0.8 , σ = 0.036515 )
P ( X > 0.86 ) = 1 - P ( X < 0.86 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 0.86 - 0.8 ) / 0.036515
Z = 1.6432
P ( ( X - µ ) / σ ) > ( 0.86 - 0.8 ) / 0.036515 )
P ( Z > 1.6432 )
P ( X > 0.86 ) = 1 - P ( Z < 1.6432 )
P ( X > 0.86 ) = 1 - 0.9498
P ( X > 0.86 ) = 0.0502